OpenStudy (gabylovesyou):

In a certain county, the number of Charter Schools is eight less than twice the number of Alternative Schools. We know that there are 36 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (anonymous):

CS: 2x-8=36 AS: 2x so is you find 2x you find the number of Alternative Schools (AS) 2x-8=36 2x=36+8 2x=44

OpenStudy (anonymous):

Correct

OpenStudy (gabylovesyou):

thanks :)

OpenStudy (anonymous):

you welcome